# A Survey of Protocol Fuzzing

### v2

@@ Acknowledgements.
- Roitburd et al. (2021) Gerbert Roitburd, Matthias Ortmann, Matthias Hollick, and Jiska Classen. 2021. Very Pwnable Network: Cisco AnyConnect Security Analysis. In 2021 IEEE Conference on Communications and Network Security (CNS). 56–64. https://doi.org/10.1109/CNS53000.2021.9705023
- Romero and Rivas (2019) Daniel Romero and Mario Rivas. 2019. Why you should fear your ’mundane’ office equipment. https://media.defcon.org/DEF%20CON%2027/DEF%20CON%2027%20presentations/DEFCON-27-Daniel-Romero-and-Mario-Rivas-Why-you-should-fear-your-mundane-office.pdf
- Rossow (2014) Christian Rossow. 2014. Amplification Hell: Revisiting Network Protocols for DDoS Abuse.. In NDSS. 1–15.
- Ruge et al. (2020) Jan Ruge, Jiska Classen, Francesco Gringoli, and Matthias Hollick. 2020. Frankenstein: Advanced Wireless Fuzzing to Exploit New Bluetooth Escalation Targets. In 29th USENIX Security Symposium (USENIX Security 20). USENIX Association, 19–36. https://www.usenix.org/conference/usenixsecurity20/presentation/ruge
- Sagonas and Typaldos (2023) Konstantinos Sagonas and Thanasis Typaldos. 2023. EDHOC-Fuzzer: An EDHOC Protocol State Fuzzer. In Proceedings of the 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis. 1495–1498.
- Sasnauskas et al. (2008) Raimondas Sasnauskas, Jó Ágila Bitsch Link, Muhammad Hamad Alizai, and Klaus Wehrle. 2008. KleeNet: Automatic Bug Hunting in Sensor Network Applications. In Proceedings of the 6th ACM Conference on Embedded Network Sensor Systems (Raleigh, NC, USA) (SenSys ’08). Association for Computing Machinery, New York, NY, USA, 425–426. https://doi.org/10.1145/1460412.1460485
- Schepers et al. (2021) Domien Schepers, Mathy Vanhoef, and Aanjhan Ranganathan. 2021. A Framework to Test and Fuzz Wi-Fi Devices. In Proceedings of the 14th ACM Conference on Security and Privacy in Wireless and Mobile Networks (Abu Dhabi, United Arab Emirates) (WiSec ’21). Association for Computing Machinery, New York, NY, USA, 368–370. https://doi.org/10.1145/3448300.3468261
- Schumilo et al. (2022) Sergej Schumilo, Cornelius Aschermann, Andrea Jemmett, Ali Abbasi, and Thorsten Holz. 2022. Nyx-Net: Network Fuzzing with Incremental Snapshots. In Proceedings of the Seventeenth European Conference on Computer Systems (Rennes, France) (EuroSys ’22). Association for Computing Machinery, New York, NY, USA, 166–180. https://doi.org/10.1145/3492321.3519591
@@
- Serebryany and Iskhodzhanov (2009) Konstantin Serebryany and Timur Iskhodzhanov. 2009. ThreadSanitizer: data race detection in practice. In Proceedings of the workshop on binary instrumentation and applications. 62–71.
@@
- Shi et al. (2019) Shangcheng Shi, Xianbo Wang, and Wing Cheong Lau. 2019. MoSSOT: An Automated Blackbox Tester for Single Sign-On Vulnerabilities in Mobile Applications. In Proceedings of the 2019 ACM Asia Conference on Computer and Communications Security (Auckland, New Zealand) (Asia CCS ’19). Association for Computing Machinery, New York, NY, USA, 269–282. https://doi.org/10.1145/3321705.3329801
@@
- Sommer et al. (2017) Manuel Sommer, Nicholas Gray, Phuoc Tran-Gia, and Thomas Zinner. 2017. FlowFuzz: A Framework for Fuzzing OpenFlow-enabled Software and Hardware Switches. https://www.blackhat.com/us-17/briefings/schedule/#flowfuzz---a-framework-for-fuzzing-openflow-enabled-software-and-hardware-switches-7642
@@
- Sommer et al. (2018) Manuel Sommer, Nicholas Gray, Phuoc Tran-Gia, and Thomas Zinner. 2018. Designing and Applying Extensible RF Fuzzing Tools to Expose PHY Layer Vulnerabilities. https://media.defcon.org/DEF%20CON%2026/DEF%20CON%2026%20presentations/DEFCON-26-Matt-Knight-and-Ryan-Speers-Designing-RF-Fuzzing-Tools-to-Expose-PHY-Layer-Vulns-Updated.pdf
@@
- Somorovsky (2016) Juraj Somorovsky. 2016. Systematic Fuzzing and Testing of TLS Libraries. In Proceedings of the 2016 ACM SIGSAC Conference on Computer and Communications Security (Vienna, Austria) (CCS ’16). Association for Computing Machinery, New York, NY, USA, 1492–1504. https://doi.org/10.1145/2976749.2978411
@@
- Song et al. (2019) Dokyung Song, Julian Lettner, Prabhu Rajasekaran, Yeoul Na, Stijn Volckaert, Per Larsen, and Michael Franz. 2019. SoK: Sanitizing for security. In 2019 IEEE Symposium on Security and Privacy (SP). IEEE, 1275–1295.
@@
- Song et al. (2014) JaeSeung Song, Cristian Cadar, and Peter Pietzuch. 2014. SymbexNet: Testing Network Protocol Implementations with Symbolic Execution and Rule-Based Specifications. IEEE Transactions on Software Engineering 40, 7 (2014), 695–709. https://doi.org/10.1109/TSE.2014.2323977
@@
- Song et al. (2022) Jonghyuk Song, Soohwan Oh, and Woongjo Choi. 2022. Automotive Ethernet Fuzzing: From Purchasing ECU to SOME/IP Fuzzing. https://forum.defcon.org/node/242347
@@
- Stepanov and Serebryany (2015) Evgeniy Stepanov and Konstantin Serebryany. 2015. MemorySanitizer: fast detector of uninitialized memory use in C++. In 2015 IEEE/ACM International Symposium on Code Generation and Optimization (CGO). IEEE, 46–55.
@@
- Stephens et al. (2016) Nick Stephens, John Grosen, Christopher Salls, Andrew Dutcher, Ruoyu Wang, Jacopo Corbetta, Yan Shoshitaishvili, Christopher Kruegel, and Giovanni Vigna. 2016. Driller: Augmenting Fuzzing Through Selective Symbolic Execution. In 23rd Annual Network and Distributed System Security Symposium, NDSS 2016, San Diego, California, USA, February 21-24, 2016. The Internet Society. http://wp.internetsociety.org/ndss/wp-content/uploads/sites/25/2017/09/driller-augmenting-fuzzing-through-selective-symbolic-execution.pdf
- Stone et al. (2021) Chris McMahon Stone, Sam L. Thomas, Mathy Vanhoef, James Henderson, Nicolas Bailluet, and Tom Chothia. 2021. The Closer You Look, The More You Learn: A Grey-box Approach to Protocol State Machine Learning. CoRR abs/2106.02623 (2021). arXiv:2106.02623 https://arxiv.org/abs/2106.02623
@@
- Sudhodanan et al. (2016) Avinash Sudhodanan, Alessandro Armando, Roberto Carbone, Luca Compagna, et al. 2016. Attack Patterns for Black-Box Security Testing of Multi-Party Web Applications.. In NDSS.
@@
- Sun et al. (2017) Wei Sun, Lisong Xu, and Sebastian Elbaum. 2017. Improving the Cost-Effectiveness of Symbolic Testing Techniques for Transport Protocol Implementations under Packet Dynamics. In Proceedings of the 26th ACM SIGSOFT International Symposium on Software Testing and Analysis (Santa Barbara, CA, USA) (ISSTA 2017). Association for Computing Machinery, New York, NY, USA, 79–89. https://doi.org/10.1145/3092703.3092706
@@
- Sun et al. (2018) Wei Sun, Lisong Xu, and Sebastian Elbaum. 2018. Scalably Testing Congestion Control Algorithms of Real-World TCP Implementations. In 2018 IEEE International Conference on Communications (ICC). 1–7. https://doi.org/10.1109/ICC.2018.8422949
@@
- Sun et al. (2019) Wei Sun, Lisong Xu, Sebastian Elbaum, and Di Zhao. 2019. Model-Agnostic and Efficient Exploration of Numerical State Space of Real-World TCP Congestion Control Implementations. In 16th USENIX Symposium on Networked Systems Design and Implementation (NSDI 19). USENIX Association, Boston, MA, 719–734. https://www.usenix.org/conference/nsdi19/presentation/sun
- Sun (2005) Zhili Sun. 2005. Satellite networking: Principles and protocols. John Wiley & Sons.
@@
- Trimberger and Moore (2014) Stephen M Trimberger and Jason J Moore. 2014. FPGA security: Motivations, features, and applications. Proc. IEEE 102, 8 (2014), 1248–1265.
@@
- Tsankov et al. (2013) Petar Tsankov, Mohammad Torabi Dashti, and David Basin. 2013. Semi-Valid Input Coverage for Fuzz Testing. In Proceedings of the 2013 International Symposium on Software Testing and Analysis (Lugano, Switzerland) (ISSTA 2013). Association for Computing Machinery, New York, NY, USA, 56–66. https://doi.org/10.1145/2483760.2483787
@@
- Vanhoef (2017) Mathy Vanhoef. 2017. WiFuzz: Detecting and Exploiting Logical Flaws in the Wi-Fi Cryptographic Handshake. https://www.blackhat.com/us-17/briefings/schedule/#wifuzz-detecting-and-exploiting-logical-flaws-in-the-wi-fi-cryptographic-handshake-6827
@@
- Walz and Sikora (2017) Andreas Walz and Axel Sikora. 2017. Exploiting dissent: towards fuzzing-based differential black-box testing of TLS implementations. IEEE Transactions on Dependable and Secure Computing 17, 2 (2017), 278–291.
- Walz and Sikora (2020) Andreas Walz and Axel Sikora. 2020. Exploiting Dissent: Towards Fuzzing-Based Differential Black-Box Testing of TLS Implementations. IEEE Transactions on Dependable and Secure Computing 17, 2 (2020), 278–291. https://doi.org/10.1109/TDSC.2017.2763947
@@
- Wang et al. (2017) Junjie Wang, Bihuan Chen, Lei Wei, and Yang Liu. 2017. Skyfire: Data-driven seed generation for fuzzing. In 2017 IEEE Symposium on Security and Privacy (SP). IEEE, 579–594.
@@
- Wang et al. (2021) Qinying Wang, Shouling Ji, Yuan Tian, Xuhong Zhang, Binbin Zhao, Yuhong Kan, Zhaowei Lin, Changting Lin, Shuiguang Deng, Alex X. Liu, and Raheem Beyah. 2021. MPInspector: A Systematic and Automatic Approach for Evaluating the Security of IoT Messaging Protocols. In 30th USENIX Security Symposium, USENIX Security 2021, August 11-13, 2021, Michael Bailey and Rachel Greenstadt (Eds.). USENIX Association, 4205–4222. https://www.usenix.org/conference/usenixsecurity21/presentation/wang-qinying
@@
- Wang and Wang (2023) Zhuzhu Wang and Ying Wang. 2023. NLP-based Cross-Layer 5G Vulnerabilities Detection via Fuzzing Generated Run-Time Profiling. arXiv preprint arXiv:2305.08226 (2023).
@@
- Wu and Li (2021) Huiyu Wu and Yuxiang Li. 2021. X-in-the-Middle: Attacking Fast Charging Piles and Electric Vehicles. https://www.blackhat.com/asia-21/briefings/schedule/#x-in-the-middle-attacking-fast-charging-piles-and-electric-vehicles--22055
@@
- Wu et al. (2021) Jianliang Wu, Ruoyu Wu, Daniele Antonioli, Mathias Payer, Nils Ole Tippenhauer, Dongyan Xu, Dave (Jing) Tian, and Antonio Bianchi. 2021. LIGHTBLUE: Automatic Profile-Aware Debloating of Bluetooth Stacks. In 30th USENIX Security Symposium (USENIX Security 21). USENIX Association, 339–356. https://www.usenix.org/conference/usenixsecurity21/presentation/wu-jianliang
@@
- Xie et al. (2020) Haikuo Xie, Ying Wang, and Ye Zhang. 2020. WIFI-Important Remote Attack Surface: Threat is Expanding. https://www.blackhat.com/asia-20/briefings/schedule/#wifi-important-remote-attack-surface-threat-is-expanding-18784
@@
- Yan et al. (2022) Han Yan, Lewei Qu, and Dongxiang Ke. 2022. BrokenMesh: New Attack Surfaces of Bluetooth Mesh. https://www.blackhat.com/us-22/briefings/schedule/#brokenmesh-new-attack-surfaces-of-bluetooth-mesh-26853
@@
- Yang et al. (2021) Youngseok Yang, Taesoo Kim, and Byung-Gon Chun. 2021. Finding Consensus Bugs in Ethereum via Multi-transaction Differential Fuzzing. In 15th USENIX Symposium on Operating Systems Design and Implementation (OSDI 21). USENIX Association, 349–365. https://www.usenix.org/conference/osdi21/presentation/yang
@@
- Yasin et al. (2020) Affan Yasin, Rubia Fatima, Lijie Wen, Wasif Afzal, Muhammad Azhar, and Richard Torkar. 2020. On Using Grey Literature and Google Scholar in Systematic Literature Reviews in Software Engineering. IEEE Access 8 (2020), 36226–36243. https://doi.org/10.1109/ACCESS.2020.2971712
@@
- Yen et al. (2021) Ta-Lun Yen, Federico Maggi, Erik Boasson, Victor Mayoral-Vilches, Mars Cheng, Patrick Kuo, and Chizuru Toyama. 2021. The Data Distribution Service (DDS) Protocol is Critical Let’s Use it Securely! https://www.blackhat.com/eu-21/briefings/schedule/#the-data-distribution-service-dds-protocol-is-critical-lets-use-it-securely-24934
@@
- Yu et al. (2019) Bo Yu, Pengfei Wang, Tai Yue, and Yong Tang. 2019. Poster: Fuzzing IoT Firmware via Multi-Stage Message Generation. In Proceedings of the 2019 ACM SIGSAC Conference on Computer and Communications Security (London, United Kingdom) (CCS ’19). Association for Computing Machinery, New York, NY, USA, 2525–2527. https://doi.org/10.1145/3319535.3363247
@@
- Yu et al. (2022) Zhenhua Yu, Haolu Wang, Dan Wang, Zhiwu Li, and Houbing Song. 2022. CGFuzzer: A Fuzzing Approach Based on Coverage-Guided Generative Adversarial Networks for Industrial IoT Protocols. IEEE Internet of Things Journal 9, 21 (2022), 21607–21619. https://doi.org/10.1109/JIOT.2022.3183952
@@
- Zhang et al. (2023a) Cen Zhang, Mingqiang Bai, Yaowen Zheng, Yeting Li, Xiaofei Xie, Yuekang Li, Wei Ma, Limin Sun, and Yang Liu. 2023a. Understanding large language model based fuzz driver generation. arXiv preprint arXiv:2307.12469 (2023).
- Zhang et al. (2021a) Cen Zhang, Yuekang Li, Hongxu Chen, Xiaoxing Luo, Miaohua Li, Anh Quynh Nguyen, and Yang Liu. 2021a. BIFF: PRactical binary fuzzing framework for programs of IoT and mobile devices. In 2021 36th IEEE/ACM International Conference on Automated Software Engineering (ASE). IEEE, 1161–1165.
@@
- Zhang et al. (2023c) Cen Zhang, Yuekang Li, Hao Zhou, Xiaohan Zhang, Yaowen Zheng, Xian Zhan, Xiaofei Xie, Xiapu Luo, Xinghua Li, Yang Liu 0003, and Sheikh Mahbub Habib. 2023c. Automata-Guided Control-Flow-Sensitive Fuzz Driver Generation. In 32nd USENIX Security Symposium, USENIX Security 2023, Anaheim, CA, USA, August 9-11, 2023, Joseph A. Calandrino and Carmela Troncoso (Eds.). USENIX Association, 2867–2884. https://www.usenix.org/conference/usenixsecurity23/presentation/zhang-cen
@@
- Zhang et al. (2021b) Cen Zhang, Xingwei Lin, Yuekang Li, Yinxing Xue, Jundong Xie, Hongxu Chen, Xinlei Ying, Jiashui Wang, and Yang Liu. 2021b. APICraft: Fuzz Driver Generation for Closed-source SDK Libraries. In 30th USENIX Security Symposium (USENIX Security 21). 2811–2828.
@@
- Zhang et al. (2023b) Zenong Zhang, George Klees, Eric Wang, Michael Hicks, and Shiyi Wei. 2023b. Fuzzing Configurations of Program Options. ACM Trans. Softw. Eng. Methodol. 32, 2, Article 53 (mar 2023), 21 pages. https://doi.org/10.1145/3580597
@@
- Zhao et al. (2019) Hui Zhao, Zhihui Li, Hansheng Wei, Jianqi Shi, and Yanhong Huang. 2019. SeqFuzzer: An Industrial Protocol Fuzzing Framework from a Deep Learning Perspective. In 12th IEEE Conference on Software Testing, Validation and Verification, ICST 2019, Xi’an, China, April 22-27, 2019. IEEE, 59–67. https://doi.org/10.1109/ICST.2019.00016
@@
- Zheng et al. (2022) Yaowen Zheng, Yuekang Li, Cen Zhang, Hongsong Zhu, Yang Liu, and Limin Sun. 2022. Efficient greybox fuzzing of applications in Linux-based IoT devices via enhanced user-mode emulation. In Proceedings of the 31st ACM SIGSOFT International Symposium on Software Testing and Analysis. 417–428.
@@
- Zheng and Sun (2022) Yaowen Zheng and Limin Sun. 2022. IPSpex: Enabling Efficient Fuzzing via Specification Extraction on ICS Protocol. In Applied Cryptography and Network Security: 20th International Conference, ACNS 2022, Rome, Italy, June 20–23, 2022, Proceedings, Vol. 13269. Springer Nature, 356.

### v3

@@ Acknowledgements.
- Roitburd et al. (2021) Gerbert Roitburd, Matthias Ortmann, Matthias Hollick, and Jiska Classen. 2021. Very Pwnable Network: Cisco AnyConnect Security Analysis. In 2021 IEEE Conference on Communications and Network Security (CNS). 56–64.
- Romero and Rivas (2019) Daniel Romero and Mario Rivas. 2019. Why you should fear your ’mundane’ office equipment. In DEFCON 27.
- Rossow (2014) Christian Rossow. 2014. Amplification Hell: Revisiting Network Protocols for DDoS Abuse. In 21st Annual Network and Distributed System Security Symposium (NDSS). 1–15.
- Ruge et al. ([n. d.]) Jan Ruge, Jiska Classen, Francesco Gringoli, and Matthias Hollick. [n. d.]. Frankenstein: Advanced Wireless Fuzzing to Exploit New Bluetooth Escalation Targets. In 29th USENIX Security Symposium (USENIX Security). 19–36.
- Sagonas and Typaldos (2023) Konstantinos Sagonas and Thanasis Typaldos. 2023. EDHOC-Fuzzer: An EDHOC Protocol State Fuzzer. In 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA). 1495–1498.
- Sasnauskas et al. (2008) Raimondas Sasnauskas, Jó Ágila Bitsch Link, Muhammad Hamad Alizai, and Klaus Wehrle. 2008. KleeNet: Automatic Bug Hunting in Sensor Network Applications. In 6th ACM Conference on Embedded Network Sensor Systems (SenSys). 425–426.
- Schepers et al. (2021) Domien Schepers, Mathy Vanhoef, and Aanjhan Ranganathan. 2021. A Framework to Test and Fuzz Wi-Fi Devices. In 14th ACM Conference on Security and Privacy in Wireless and Mobile Networks (WiSec). 368–370.
- Schumilo et al. (2022) Sergej Schumilo, Cornelius Aschermann, Andrea Jemmett, Ali Abbasi, and Thorsten Holz. 2022. Nyx-Net: Network Fuzzing with Incremental Snapshots. In 17th European Conference on Computer Systems (EuroSys). 166–180.
@@
- Serebryany and Iskhodzhanov (2009) Konstantin Serebryany and Timur Iskhodzhanov. 2009. ThreadSanitizer: data race detection in practice. In workshop on binary instrumentation and applications. 62–71.
@@
- Shelby (2014) Z. Shelby. 2014. RFC7252: The Constrained Application Protocol (CoAP). https://www.rfc-editor.org/rfc/rfc7252
- Shi et al. (2023a) Qingkai Shi, Junyang Shao, Yapeng Ye, Mingwei Zheng, and Xiangyu Zhang. 2023a. Lifting Network Protocol Implementation to Precise Format Specification with Security Applications. In 2023 ACM SIGSAC Conference on Computer and Communications Security (CCS). 1287–1301.
- Shi et al. (2023b) Qingkai Shi, Xiangzhe Xu, and Xiangyu Zhang. 2023b. Extracting protocol format as state machine via controlled static loop analysis. In 32nd USENIX Security Symposium (USENIX Security 23). 7019–7036.
- Shi et al. (2019) Shangcheng Shi, Xianbo Wang, and Wing Cheong Lau. 2019. MoSSOT: An Automated Blackbox Tester for Single Sign-On Vulnerabilities in Mobile Applications. In 2019 ACM Asia Conference on Computer and Communications Security (AsiaCCS). 269–282.
@@
- Sommer et al. (2017) Manuel Sommer, Nicholas Gray, Phuoc Tran-Gia, and Thomas Zinner. 2017. FlowFuzz: A Framework for Fuzzing OpenFlow-enabled Software and Hardware Switches. In Blackhat US 2017.
@@
- Sommer et al. (2018) Manuel Sommer, Nicholas Gray, Phuoc Tran-Gia, and Thomas Zinner. 2018. Designing and Applying Extensible RF Fuzzing Tools to Expose PHY Layer Vulnerabilities. In DEFCON 26.
@@
- Somorovsky (2016) Juraj Somorovsky. 2016. Systematic Fuzzing and Testing of TLS Libraries. In 2016 ACM SIGSAC Conference on Computer and Communications Security (CCS). 1492–1504.
@@
- Song et al. (2019) Dokyung Song, Julian Lettner, Prabhu Rajasekaran, Yeoul Na, Stijn Volckaert, Per Larsen, and Michael Franz. 2019. SoK: Sanitizing for security. In 2019 IEEE Symposium on Security and Privacy (SP). 1275–1295.
@@
- Song et al. (2014) JaeSeung Song, Cristian Cadar, and Peter Pietzuch. 2014. SymbexNet: Testing Network Protocol Implementations with Symbolic Execution and Rule-Based Specifications. IEEE Transactions on Software Engineering 40, 7 (2014), 695–709.
@@
- Song et al. (2022) Jonghyuk Song, Soohwan Oh, and Woongjo Choi. 2022. Automotive Ethernet Fuzzing: From Purchasing ECU to SOME/IP Fuzzing. In DEFCON 30.
@@
- Stepanov and Serebryany (2015) Evgeniy Stepanov and Konstantin Serebryany. 2015. MemorySanitizer: fast detector of uninitialized memory use in C++. In 2015 IEEE/ACM International Symposium on Code Generation and Optimization (CGO). 46–55.
@@
- Stephens et al. (2016) Nick Stephens, John Grosen, Christopher Salls, Andrew Dutcher, Ruoyu Wang, Jacopo Corbetta, Yan Shoshitaishvili, Christopher Kruegel, and Giovanni Vigna. 2016. Driller: Augmenting Fuzzing Through Selective Symbolic Execution. In 23rd Annual Network and Distributed System Security Symposium (NDSS).
@@
- Sudhodanan et al. (2016) Avinash Sudhodanan, Alessandro Armando, Roberto Carbone, Luca Compagna, et al. 2016. Attack Patterns for Black-Box Security Testing of Multi-Party Web Applications.. In 23st Annual Network and Distributed System Security Symposium (NDSS).
@@
- Sun et al. (2017) Wei Sun, Lisong Xu, and Sebastian Elbaum. 2017. Improving the Cost-Effectiveness of Symbolic Testing Techniques for Transport Protocol Implementations under Packet Dynamics. In 26th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA). 79–89.
@@
- Sun et al. (2018) Wei Sun, Lisong Xu, and Sebastian Elbaum. 2018. Scalably Testing Congestion Control Algorithms of Real-World TCP Implementations. In 2018 IEEE International Conference on Communications (ICC). 1–7.
@@
- Sun et al. (2019) Wei Sun, Lisong Xu, Sebastian Elbaum, and Di Zhao. 2019. Model-Agnostic and Efficient Exploration of Numerical State Space of Real-World TCP Congestion Control Implementations. In 16th USENIX Symposium on Networked Systems Design and Implementation (NSDI 19). 719–734.
- Sun et al. (2023) Yue Sun, Zhi Li, Shichao Lv, and Limin Sun. 2023. Spenny: Extensive ICS Protocol Reverse Analysis via Field Guided Symbolic Execution. IEEE Transactions on Dependable and Secure Computing 20, 6 (2023), 4502–4518.
- Sun (2014) Zhili Sun. 2014. Satellite Networking: Principles and Protocols (2nd ed.). Wiley Publishing.
@@
- Trimberger and Moore (2014) Stephen M Trimberger and Jason J Moore. 2014. FPGA security: Motivations, features, and applications. IEEE 102, 8 (2014), 1248–1265.
@@
- Tsankov et al. (2013) Petar Tsankov, Mohammad Torabi Dashti, and David Basin. 2013. Semi-Valid Input Coverage for Fuzz Testing. In 2013 International Symposium on Software Testing and Analysis (ISSTA). 56–66.
@@
- Vanhoef (2017) Mathy Vanhoef. 2017. WiFuzz: Detecting and Exploiting Logical Flaws in the Wi-Fi Cryptographic Handshake. In Blackhat US 2017.
@@
- Walz and Sikora (2020) Andreas Walz and Axel Sikora. 2020. Exploiting Dissent: Towards Fuzzing-Based Differential Black-Box Testing of TLS Implementations. IEEE Transactions on Dependable and Secure Computing 17, 2 (2020), 278–291.
@@
- Wang et al. (2017) Junjie Wang, Bihuan Chen, Lei Wei, and Yang Liu. 2017. Skyfire: Data-driven seed generation for fuzzing. In 2017 IEEE Symposium on Security and Privacy (SP). 579–594.
- Wang et al. (2024) J. Wang, L. Yu, and X. Luo. 2024. LLMIF: Augmented Large Language Model for Fuzzing IoT Devices. In 2024 IEEE Symposium on Security and Privacy (SP). 196–196.
@@
- Wang et al. (2021) Qinying Wang, Shouling Ji, Yuan Tian, Xuhong Zhang, Binbin Zhao, Yuhong Kan, Zhaowei Lin, Changting Lin, Shuiguang Deng, Alex X. Liu, and Raheem Beyah. 2021. MPInspector: A Systematic and Automatic Approach for Evaluating the Security of IoT Messaging Protocols. In 30th USENIX Security Symposium (USENIX Security). 4205–4222.
@@
- Wang and Wang (2023) Zhuzhu Wang and Ying Wang. 2023. NLP-based Cross-Layer 5G Vulnerabilities Detection via Fuzzing Generated Run-Time Profiling. arXiv:2305.08226 (2023).
- Wu et al. (2024) Feifan Wu, Zhengxiong Luo, Yanyang Zhao, Qingpeng Du, Junze Yu, Ruikang Peng, Heyuan Shi, and Yu Jiang. 2024. Logos: Log Guided Fuzzing for Protocol Implementations. In 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA). 1–13.
@@
- Wu and Li (2021) Huiyu Wu and Yuxiang Li. 2021. X-in-the-Middle: Attacking Fast Charging Piles and Electric Vehicles. In Blackhat Asia 2021.
@@
- Wu et al. (2021) Jianliang Wu, Ruoyu Wu, Daniele Antonioli, Mathias Payer, Nils Ole Tippenhauer, Dongyan Xu, Dave (Jing) Tian, and Antonio Bianchi. 2021. LIGHTBLUE: Automatic Profile-Aware Debloating of Bluetooth Stacks. In 30th USENIX Security Symposium (USENIX Security 21). 339–356.
@@
- Xie et al. (2020) Haikuo Xie, Ying Wang, and Ye Zhang. 2020. WIFI-Important Remote Attack Surface: Threat is Expanding. In Blackhat Asia 2020.
@@
- Yan et al. (2022) Han Yan, Lewei Qu, and Dongxiang Ke. 2022. BrokenMesh: New Attack Surfaces of Bluetooth Mesh. In Blackhat US 2022.
@@
- Yang et al. (2021) Youngseok Yang, Taesoo Kim, and Byung-Gon Chun. 2021. Finding Consensus Bugs in Ethereum via Multi-transaction Differential Fuzzing. In 15th USENIX Symposium on Operating Systems Design and Implementation (OSDI). 349–365.
@@
- Yasin et al. (2020) Affan Yasin, Rubia Fatima, Lijie Wen, Wasif Afzal, Muhammad Azhar, and Richard Torkar. 2020. On Using Grey Literature and Google Scholar in Systematic Literature Reviews in Software Engineering. IEEE Access 8 (2020), 36226–36243.
- Ye et al. (2023) Mingxi Ye, Yuhong Nan, Zibin Zheng, Dongpeng Wu, and Huizhong Li. 2023. Detecting state inconsistency bugs in dapps via on-chain transaction replay and fuzzing. In 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA). 298–309.
@@
- Yen et al. (2021) Ta-Lun Yen, Federico Maggi, Erik Boasson, Victor Mayoral-Vilches, Mars Cheng, Patrick Kuo, and Chizuru Toyama. 2021. The Data Distribution Service (DDS) Protocol is Critical Let’s Use it Securely!. In Blackhat EU 2021.
@@
- Yu et al. (2019) Bo Yu, Pengfei Wang, Tai Yue, and Yong Tang. 2019. Poster: Fuzzing IoT Firmware via Multi-Stage Message Generation. In 2019 ACM SIGSAC Conference on Computer and Communications Security (CCS). 2525–2527.
- Yu et al. (2023) Le Yu, Rui Yao, and Zhanlei Zhang. 2023. Poster: DP-Reverser: Automatically Reverse Engineering Vehicle Diagnostic Protocols. In 2023 IEEE 43rd International Conference on Distributed Computing Systems (ICDCS). 1053–1054.
@@
- Yu et al. (2022) Zhenhua Yu, Haolu Wang, Dan Wang, Zhiwu Li, and Houbing Song. 2022. CGFuzzer: A Fuzzing Approach Based on Coverage-Guided Generative Adversarial Networks for Industrial IoT Protocols. IEEE Internet Things J. 9, 21 (2022), 21607–21619.
@@
- Zhang et al. (2023b) Cen Zhang, Yuekang Li, Hao Zhou, Xiaohan Zhang, Yaowen Zheng, Xian Zhan, Xiaofei Xie, Xiapu Luo, Xinghua Li, Yang Liu, et al. 2023b. Automata-Guided Control-Flow-Sensitive Fuzz Driver Generation.. In 32nd USENIX Security Symposium. 2867–2884.
@@
- Zhang et al. (2021) Cen Zhang, Xingwei Lin, Yuekang Li, Yinxing Xue, Jundong Xie, Hongxu Chen, Xinlei Ying, Jiashui Wang, and Yang Liu. 2021. APICraft: Fuzz driver generation for closed-source SDK libraries. In 30th USENIX Security Symposium (USENIX Security 21). 2811–2828.
- Zhang et al. (2024) Cen Zhang, Yaowen Zheng, Mingqiang Bai, Yeting Li, Wei Ma, Xiaofei Xie, Yuekang Li, Limin Sun, and Yang Liu. 2024. How Effective Are They? Exploring Large Language Model Based Fuzz Driver Generation. In 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA). 1–13.
@@
- Zhang et al. (2023a) Zenong Zhang, George Klees, Eric Wang, Michael Hicks, and Shiyi Wei. 2023a. Fuzzing Configurations of Program Options. ACM Trans. Softw. Eng. Methodol. 32, 2 (2023), 1–21.
@@
- Zhao et al. (2019) Hui Zhao, Zhihui Li, Hansheng Wei, Jianqi Shi, and Yanhong Huang. 2019. SeqFuzzer: An Industrial Protocol Fuzzing Framework from a Deep Learning Perspective. In 12th IEEE Conference on Software Testing, Validation and Verification (ICST). 59–67.
- Zheng et al. (2024) Mingwei Zheng, Qingkai Shi, Xuwei Liu, Xiangzhe Xu, Le Yu, Congyu Liu, Guannan Wei, and Xiangyu Zhang. 2024. ParDiff: Practical Static Differential Analysis of Network Protocol Parsers. Proceedings of the ACM on Programming Languages (OOPSLA) 8 (2024), 1208–1234.
@@
- Zheng et al. (2022) Yaowen Zheng, Yuekang Li, Cen Zhang, Hongsong Zhu, Yang Liu, and Limin Sun. 2022. Efficient greybox fuzzing of applications in Linux-based IoT devices via enhanced user-mode emulation. In 31st ACM SIGSOFT International Symposium on Software Testing and Analysis. 417–428.
@@
- Zheng and Sun (2022) Yaowen Zheng and Limin Sun. 2022. IPSpex: Enabling Efficient Fuzzing via Specification Extraction on ICS Protocol. In Applied Cryptography and Network Security: 20th International Conference (ACNS), Vol. 13269. 356.